need help..The altitude in feet of a projectile t seconds after launch is given by the function s(t)=-16t^2+32t+3. How long after launch will it take for the projectile to reach its maximum height? What is the maximum height?
thanks for your help

Answer by stanbon(48564)

s(t)=-16t^2+32t+3. How long after launch will it take for the projectile to reach its maximum height? What is the maximum height?
---------------
You have a quadratic with a=-16, b=32
Since a is negative, the function has a maximum at x=-b/(2a)
=-32/(-32)=1 second.
The height at that time is:
f(1)=-16+32+3=19 ft.
Cheers,
Stan H.